OpenStudy (abdumu16):

Estimate the indicated derivative by any method. s = t − t^2; estimate ds/dt t = 5

OpenStudy (abdumu16):

@sleepyjess and @ganeshie8 and @mathmate can you guys help me out?

OpenStudy (jiteshmeghwal9):

\[\frac{ds}{dt}=\frac{dt}{dt}-\frac{dt^2}{dt}\]

OpenStudy (jiteshmeghwal9):

After getting the derivative replace t by 5

OpenStudy (mathmale):

I assume you know the concept of derivative and can apply it. Find the derivative of s with respect to time, and then let t=5 sec. Please show your work.

OpenStudy (trojanpoem):

ds/st = 1 - 2t = 1- 10 = -9

OpenStudy (mathmale):

Another way: Graph the function s = t - t^2 for t= 3, 4, 5, 6, 7. From your graph, read off the approx. slope of the tangent line to this curve at t=5.